50 km/hr = ___ m/s 14.3 m/s = ___ cm/s 5000 m/s
Answer #1

4)

a) Area time distance

A = m^2 and d = m

A x d = m^2 x m = m^3 (Area time distance is Volume)

b)Energy time acceleration:

Energy = J and acc = m/s^2

E x a = J m/s^2

c)distance times velocity

distance = m and velcoty = m/s

d x v = m^2/s

d)distance times volume

distance = m ; volume = m^3

d x v = m^4

e)Area / distance

Area = m^2 ; distance = m

Area/dist = m^2/m = m

f)Velocity/acc

Velocity = m/s and acc = m/s^2

v/a = m/s / m/s^2 = s (its time)

g)Energy divided by acc

Energy is J = N m = kg m/s^2 x m; acc = m/s^2

E/acc = kg-m/s^2 x m / m/s^2 = kg-m

energy/acc = kg/m